# V717 Andromedae: An active low mass ratio contact binary

### v1

@@ 1. Introduction
Close binary pairs where both components have expanded beyond their respective Roche lobes, resulting in bidirectional mass transfer and the formation of shared outer photospheric envelope, are referred to as contact binaries. Both components are still on the main sequence, but have expanded beyond their Roche lobes due to their close proximity and tidal effects. Although examples of high mass contact binaries are known, most examples have a primary with spectral class K to F with secondary component smaller than the primary. Regardless of the difference in the masses of the components, the temperature of the common envelope is close to that of the primary component. The secondary therefore is normally significantly brighter than its main sequence counterpart (Mochnacki, 1981; Yildiz & Doğan, 2013).
@@ 2. Target Selection and Photometric Observations
Evolution modelling of contact binary systems has mostly suggested that the components will eventually merge following a period of angular momentum loss and possible non-conservative mass loss from the outer Roche lobe of the secondary or both components. Modelling has usually indicated merger taking place at extreme low mass ratios (Robertson & Eggleton, 1977). Theoretical considerations for orbital stability initially concentrated on the concept of a global minimum mass ratio at which merger would take place (Rasio, 1995; Arbutina, 2007). More recently, however, this has been shown to be an over simplification, and it is now clear that each system has its own unique mass ratio at which it will merge and this mass ratio changes even during normal main sequence evolution (Wadhwa et al., 2021, 2024b). As part of our ongoing work in the identification of extreme low mass ratio, potentially unstable, contact binary systems, we have developed a rapid method of identification based on inspection of survey photometry (Wadhwa et al., 2022a). The rapid identification routine indicated V717 And as a potential extreme low mass ratio system and we present an analysis of multi-band photometry along with a review of potential increased chromospheric activity, long-term period analysis, and of orbital stability.
V717 And was observed at the Astronomical Station Vidojevica, operated by the Astronomical Observatory of Belgrade, on September 1, 2024, using the $1.4\,\mathrm{m}$ “Milanković” telescope equipped with an Andor iKon-L 936 CCD camera. Images were acquired using Johnson BVRB\,V\,R filters. All images were calibrated using appropriate flat, dark and bias frames. Differential photometry was performed using the AstroImageJ software package (Collins et al., 2017) with 2MASS00122578+3124054 as the comparison star and 2MASS00122452+3122164 as the check star. Reference magnitudes of the comparison star ($V=14.37$, $B=15.51$, $R=14.04$) were adopted from Henden et al. (2015). Basic observations of the light curves, summarised in Table 1, show a $V$-band amplitude of 0.41 mag with essentially equal maxima and minima. The $R$-band amplitude is 0.40 mag, again with near equal mimima and maxima. The $B$-band amplitude is higher at 0.48 mag, with the second maxima brighter by 0.03 mag and the primary minima fainter by about 0.03 mag. Such variations in amplitude and variation in maxima and minima between different bands are not uncommon among contact binaries (Xu et al., 2022; Samec et al., 2011).
@@ 3. Period Study
There are no dedicated historical observations for V717 And. Survey data are available from the Catalina Sky Survey (Drake et al., 2014), All Sky Automated Survey - Supernova (Shappee et al., 2014; Jayasinghe et al., 2018), and the Zwicky Transient Facility Catalog of Periodic Variable Stars (Chen et al., 2020). The system was also observed by the SuperWasp survey (Thiemann et al., 2021), but the system is at the fainter limit of the instruments resulting in considerable scatter, and so no SuperWasp data were used. In total, about 800 observations spanning 6350 days were available for review. Period changes, especially when small, are most accurately determined from long-term high-cadence observations. In the absence of such observations, the traditional Observed-Computed (O-C) technique offers no meaningful analysis. The technique of using periodic orthogonal polynomials combined with an analysis of variance statistic (a marker of quality of fit), as described by Schwarzenberg-Czerny (1996) and used by Tylenda et al. (2011) in the period analysis of survey photometry of the only confirmed contact binary merger event, offers an avenue to determine significant period changes. Using the methodology described in Schwarzenberg-Czerny (1996), we divided the available survey data into multiple overlapping subsets of approximately 100 observations and determined the best period in the small range (0.25d to 0.35d) near the current estimate of 0.317357d. We find no significant change in the period of the system over 17 or so years of available observations. The changes in period over the survey time are illustrated in Figure 1. Based on our observations and the available survey data, we refine the ephemeris as follows:
@@
Figure 1: Period of V717 And spanning over 6300 days. Black line = Line of best fit
@@ 4. Light Curve Analysis and Absolute Parameters
The light curve of V717 And demonstrates complete eclipses and is therefore suitable for photometric determination of the mass ratio and geometric parameters (Terrell & Wilson, 2005). We used a recent version of the Wilson-Devinney (WD) code (Wilson & Devinney, 1971; Wilson, 1990; Nelson, 2021) to simultaneously model the BVRB\,V\,R photometric data. The widely used grid search method was used for values of the mass ratio ($q$) from 0.1 to 10 at intervals of 0.5 to find an approximate solution. Further refinement was performed around the approximate solution at intervals of 0.1 and finally at intervals of 0.01. During iterations the temperature of the secondary component ($T_{2}$), the potential (fillout) ($f$), inclination ($i$) and the dimensionless scaling factor, the luminosity of the primary ($L_{1}$), were the adjustable parameters. The fixed temperature of the primary ($T_{1}=5895K$) was adopted from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) spectra (Luo et al., 2022). As usual, the gravity coefficients were fixed at ($g_{1}=g_{2}=0.32)$), bolometric albedoes ($A_{1}=A_{2}=0.5$) were also fixed while the limb darkening coefficients were interpolated from van Hamme (1993). Simple reflection treatment was applied during analysis.
@@
The system is found to be an extreme low mass ratio contact binary with a mass ratio $q=0.197\pm 0.003$, high inclination $i=84.5^{\circ}\pm 1.0^{\circ}$, moderate contact $f=27\pm 7\%$ and near equal component temperatures, $T_{2}=5813\pm 15K$. The geometric solution is summarised in Table 2. Observed and fitted light curves are illustrated in Figure 2. The variation in amplitude and the slight O’Connell effect noted in the $B$-band photometry is reflected in the slightly poorer fit of the observed and modelled light curves.
@@
The most critical absolute parameter for further analysis is the mass of the primary component. As has been shown previously (Wadhwa et al., 2022a), the radius of the primary component of contact binaries is significantly larger than that of the main sequence counterpart, and small variations in the assignment of the temperature can lead to significant variation in the estimate of the mass of a star if simple black-body approximations are applied. For this reason, we prefer a more direct observational approach to estimate the mass of the primary component. As the light curve shows complete eclipses, we can conclude that the apparent magnitude at phase 0.5 is the apparent magnitude of the primary component. Using an accurate Gaia (Gaia Collaboration et al., 2016, 2023) distance (862.4 pc) and distance corrected extinction (0.162 mag) as described in Wadhwa et al. (2023), we derive the absolute magnitude of the primary component, $M_{V1}=4.79\pm 0.04$. Using the April 2022 update of Pecaut & Mamajek (2013) tables we estimate the mass of the primary based on the absolute magnitude as $M_{1}=1.02M_{\odot}$, in good agreement with the spectral colour estimate from Gaia of $0.97M_{\odot}$. Based on these two estimates, we adopt the mass of the primary as $M_{1}=1.00\pm 0.03M_{\odot}$. All subsequent parameters and errors are propagated from this estimate. The mass of the secondary ($M_{2}=0.197\pm 0.004M_{\odot}$) follows from the mass ratio, and the separation ($A=2.13\pm 0.02R_{\odot}$) follows from the period and Kepler’s third law. The radii of the components can be estimated as follows: ($R_{1,2}=A\times r_{1}$ and $A\times r_{2}$) where ($r_{1,2}$) are the geometric means of the fractional radii of the components in three orientations determined by the WD analysis (Awadalla & Hanna, 2005). The absolute parameters are summarised in Table 2.
@@ 5.2 Orbital Stability
Ever since the identification of V1309 Sco as a contact binary merger event (Tylenda et al., 2011), there has been a significant increase in the identification of potential merger candidates. Theoretical considerations prior to the identification of V1309 Sco had indicated merger would likely occur when the mass ratio of the components was very small and much of the research had centered on defining a universal minimum mass ratio at which merger would take place (Rasio, 1995; Arbutina, 2007). More recent work (Wadhwa et al., 2021, 2024b) has indicated that there is no universal instability mass ratio and each contact binary has a unique mass ratio at which it will become unstable. The instability mass ratio, for a given system, is not fixed and changes significantly even during normal main sequence evolution.
@@
Based on the above relations the instability mass ratio range for V717 And would be 0.100–0.118. Wadhwa et al. (2024b) further showed that if age and metallicity are taken into account, systems with metallicity less than Solar metallicity and older systems may well be more stable. Assuming the the primary is half way through its main sequence evolution and adopting the LAMOST metallicity of $\mathrm{Fe}/\mathrm{H}]=-0.38$, the predicted instability mass ratio range is even lower at 0.065–0.060. The estimated mass ratio of V717 And at 0.197 is well above the instability mass ratio and as such we consider the system to be stable and not a potential merger candidate.
@@ 6. Summary
Although the system is of extreme low mass ratio, orbital stability analysis suggests that the system is stable and not a potential merger candidate.

### v2

@@ 1. Introduction
Close binary pairs where both components have expanded beyond their respective Roche lobes, resulting in bidirectional mass transfer and the formation of shared outer photospheric envelope, are referred to as contact binaries. Both components are still on the main sequence, but have expanded beyond their Roche lobes due to their close proximity and tidal effects. Although examples of high mass contact binaries are known, most examples have a primary (larger and heavier of the two components) with spectral class K to F with secondary component smaller and lighter than the primary. Regardless of the difference in the masses of the components, the temperature of the common envelope is close to that of the primary component. The secondary therefore is normally significantly brighter than its main sequence counterpart (Mochnacki, 1981; Yildiz & Doğan, 2013). Recently, there has been significant interest in the study of extreme low mass ratio contact binaries and theoretical considerations suggest that such systems may be unstable and potential merger candidates. Although quite a few extreme low mass ratio contact binaries have been found their predicted merger induced transient, a red nova, remains rare with only one confirmed event (Tylenda et al., 2011). The rarity of merger events suggests that pre-merger stage(s) is/are critical and poorly understood; identification of potential merger candidates therefore has gained considerable interest recently. Continued investigation of extreme low mass ratio contact binaries can potentially constrain the mechanisms underlying merger, fully define the instability mass ratio and potentially allow real time observation of stellar mergers.
@@ 2. Target Selection and Photometric Observations
Evolution modelling of contact binary systems has mostly suggested that the components will eventually merge following a period of angular momentum loss and possible non-conservative mass loss from the outer Roche lobe of the secondary or both components. Modelling has usually indicated merger taking place at extreme low mass ratios (Robertson & Eggleton, 1977). Theoretical considerations for orbital stability initially concentrated on the concept of a global minimum mass ratio at which merger would take place (Rasio, 1995; Arbutina, 2007). More recently, however, this has been shown to be an over simplification, and it is now clear that each system has its own unique mass ratio at which it will merge and this mass ratio changes even during normal main sequence evolution (Wadhwa et al., 2021, 2024b). As part of our ongoing work in the identification of extreme low mass ratio, potentially unstable, contact binary systems, we have developed a rapid method of identification based on inspection of survey photometry (Wadhwa et al., 2022b). The rapid identification routine indicated V717 And as a potential extreme low mass ratio system and we present an analysis of multi-band photometry along with a review of potential increased chromospheric activity, long-term period analysis, and of orbital stability.
V717 And was observed at the Astronomical Station Vidojevica, operated by the Astronomical Observatory of Belgrade, on September 1, 2024, using the $1.4\,\mathrm{m}$ “Milanković” telescope equipped with an Andor iKon-L 936 CCD camera. Images were acquired using Johnson BVRB\,V\,R filters. Exposure times were 90 seconds, 45 seconds and 30 seconds for BVRB\,V\,R filters respectively. All images were calibrated using appropriate flat, dark and bias frames. Differential photometry was performed using the AstroImageJ software package (Collins et al., 2017) with 2MASS00122578+3124054 as the comparison star and 2MASS00122452+3122164 as the check star. Reference magnitudes of the comparison star ($V=14.37$, $B=15.51$, $R=14.04$) were adopted from Henden et al. (2015). The mean reported error for the estimated magnitudes was less than 0.005 mag for all filters. Basic observations of the light curves, summarised in Table 1, show a $V$-band amplitude of 0.41 mag with essentially equal maxima and minima. The $R$-band amplitude is 0.40 mag, again with near equal mimima and maxima. The $B$-band amplitude is higher at 0.48 mag, with the second maxima brighter by 0.03 mag and the primary minima fainter by about 0.03 mag. Such variations in amplitude and variation in maxima and minima between different bands are not uncommon among contact binaries (Xu et al., 2022; Samec et al., 2011).
@@ 3. Period Study
There are no dedicated historical observations for V717 And. Survey data are available from the Catalina Sky Survey (Drake et al., 2014), All Sky Automated Survey - Supernova (Shappee et al., 2014; Jayasinghe et al., 2018), and the Zwicky Transient Facility Catalog of Periodic Variable Stars (Chen et al., 2020). The system was also observed by the SuperWasp survey (Thiemann et al., 2021), but the system is at the fainter limit of its instruments resulting in considerable scatter such that SuperWasp data was not used. In total, about 800 observations spanning 6350 days were available for review. Period changes, especially when small, are most accurately determined from long-term high-cadence observations. In the absence of such observations, the traditional Observed-Computed (O-C) technique offers no meaningful analysis. The technique of using periodic orthogonal polynomials to fit low cadence observations combined with an analysis of variance statistic to evaluate the quality of fit is regarded as being highly suitable for analysing unevenly and widely separated observation. The methodology is fully described in (Schwarzenberg-Czerny, 1996) and is widely implimented in various software packages including ”Peranso” (Paunzen & Vanmunster, 2016). Briefly, the data sample is divided into smaller subsets (often overlapping) and the ”best” period for each subset calculated. The technique was used by Tylenda et al. (2011) in the period analysis of survey photometry, often using very small sample subsets, of the only confirmed contact binary merger event. Using the Peranso software package we divided the available survey data into multiple overlapping subsets of approximately 100 observations and determined the best period in the small range (0.25d to 0.35d) near the current estimate of 0.317357d. We find no significant change in the period of the system over 17 or so years of available observations. The changes in period over the survey time are illustrated in Figure 1. Based on our observations and the available survey data, we refine the ephemeris as follows:
@@
Figure 1: Period of V717 And spanning over 6300 days. Black line = Line of best fit. Estimates from various surveys as per colour coded data points.
@@ 4. Light Curve Analysis and Absolute Parameters
The light curve of V717 And demonstrates complete eclipses and is therefore suitable for photometric determination of the mass ratio and geometric parameters (Terrell & Wilson, 2005). We used a recent version of the Wilson-Devinney (WD) code (Wilson & Devinney, 1971; Wilson, 1990; Nelson, 2021) to simultaneously model the BVRB\,V\,R photometric data. The widely used grid search method was used for values of the mass ratio ($q$) from 0.1 to 10 at intervals of 0.5 to find an approximate solution. Further refinement was performed around the approximate solution at intervals of 0.1 and finally at intervals of 0.01. During iterations the temperature of the secondary component ($T_{2}$), the potential (fillout) ($f$), inclination ($i$) and the dimensionless scaling factor, the luminosity of the primary ($L_{1}$), were the adjustable parameters. The fixed temperature of the primary ($T_{1}=5895K$) was adopted from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) spectra (Luo et al., 2022). As usual, the gravity coefficients were fixed at ($g_{1}=g_{2}=0.32)$), bolometric albedoes ($A_{1}=A_{2}=0.5$) were also fixed while the limb darkening coefficients were interpolated from van Hamme (1993). Simple reflection treatment was applied during analysis. As is standard practice, we adopt the standard deviation reported by the WD code on its final iteration as the errors for each estimated parameter. Potential random errors are not taken into account and as such the errors may be underestimated.
@@
The system is found to be an extreme low mass ratio contact binary with a mass ratio $q=0.197\pm 0.003$, high inclination $i=84.5^{\circ}\pm 1.0^{\circ}$, moderate contact $f=27\pm 7\%$ and near equal component temperatures, $T_{2}=5813\pm 15K$. The geometric solution is summarised in Table 2. The critical portion of the grid search is illustrated in Figure 2. Observed and fitted light curves are illustrated in Figure 3. The variation in amplitude and the slight O’Connell effect noted in the $B$-band photometry is reflected in the slightly poorer fit of the observed and modelled light curves. We found no improvement in the fit with the inclusion of a third light.
Figure 2: The critical portion of the mass ratio search grid from q=0.16 to q=0.22. The residual between the observed and fitted light curves has been normalised to the minimum residual at q=0.197
As has been noted previously (Wadhwa et al., 2021, 2022a; Devarapalli et al., 2020), analysis of survey photometric data can yield results in line with dedicated observations; however, there can be some variations, particularly with respect to the mass ratio and degree of contact. This is also true for V717 And for which analysis of survey photometry (Li et al., 2024) yields a mass ratio of 0.16, fillout of 0.43 and inclination of 80.4deg.
@@
The most critical absolute parameter for further analysis is the mass of the primary component. In the absence of spectra, the luminosity of the primary component is usually used as an indirect method to estimate the mass of the primary component. The blackbody estimate of luminosity is based on the estimated radius and temperature of the primary component. Both of these parameters have some issues, noted below, which make blackbody estimate of luminosity not the ideal choice. As has been shown previously (Wadhwa et al., 2022b), the radius of the primary component of contact binaries is significantly larger than that of the main sequence counterpart, and in the absence of formal spectra, colour based estimates of temperature can vary significantly and small variations in the assignment of the temperature can lead to significant variation in the estimate of the luminosity of a star if simple blackbody approximations are applied. For this reason, we prefer a more direct observational approach to estimate the mass of the primary component. As the light curve shows complete eclipses, we can conclude that the apparent magnitude at phase 0.5 is the apparent magnitude of the primary component. Using an accurate Gaia (Gaia Collaboration et al., 2016, 2023) distance (862.4 pc) and distance corrected extinction (0.162 mag) as described in Wadhwa et al. (2023), we derive the absolute magnitude of the primary component, $M_{V1}=4.47\pm 0.04$. Using the April 2022 update  [^1] https://tinyurl.com/7yxyry8n of Pecaut & Mamajek (2013) tables we estimate the mass of the primary based on the absolute magnitude as $M_{1}=1.04M_{\odot}$, in good agreement with the spectral colour estimate from Gaia of $0.97M_{\odot}$. Based on these two estimates, we adopt the mass of the primary as $M_{1}=1.00\pm 0.03M_{\odot}$. All subsequent parameters and errors are propagated from this estimate. The mass of the secondary ($M_{2}=0.197\pm 0.005M_{\odot}$) follows from the mass ratio, and the separation ($A=2.13\pm 0.02R_{\odot}$) follows from the period and Kepler’s third law. The radii of the components can be estimated as follows: ($R_{1,2}=A\times r_{1}$ and $A\times r_{2}$) where ($r_{1,2}$) are the geometric means of the fractional radii of the components in three orientations determined by the WD analysis (Awadalla & Hanna, 2005). The absolute parameters are summarised in Table 2.
@@ 5.2 Orbital Stability
Ever since the identification of V1309 Sco as a contact binary merger event (Tylenda et al., 2011), there has been a significant increase in the identification of potential merger candidates. Theoretical considerations prior to the identification of V1309 Sco had indicated merger would likely occur when the mass ratio of the components was very small and much of the research had centered on defining a universal minimum mass ratio at which merger would take place (Rasio, 1995; Arbutina, 2007). More recent work (Wadhwa et al., 2021, 2024b) has indicated that there is no universal instability mass ratio and each contact binary has a unique mass ratio at which it will become unstable. The instability mass ratio, for a given system, is not fixed and changes significantly even during normal main sequence evolution. Observationally there are a few indicators which point to variable and changing mass ratio of contact binaries. Pešta & Pejcha (2023) indicated that the instability mass ratio may depend on the period with extremely low instability mass ratio for systems with periods greater than 0.3d while the instability mass ratio was estimated significantly higher for systems with shorter periods. The paucity of observed red novae, even in the presence of significant and ever increasing number of extreme low mass ratio contact binaries, is also supportive of systems remaining stable potentially throughout their nuclear lifespans (Wadhwa et al., 2024b).
@@
Based on the above relations the instability mass ratio range for V717 And would be 0.100–0.118. Wadhwa et al. (2024b) further showed that if age and metallicity are taken into account, systems with metallicity less than Solar metallicity and older systems may well be more stable. Assuming the primary is half way through its main sequence evolution and adopting the LAMOST metallicity of $\mathrm{Fe}/\mathrm{H}]=-0.38$, the predicted instability mass ratio range is even lower at 0.065–0.070. The estimated mass ratio of V717 And at 0.197 is well above the instability mass ratio and as such we consider the system to be stable and not a potential merger candidate.
@@ 6. Summary
Orbital stability modelling suggests that contact binaries should merge. Theoretical considerations indicate such a merger would likely occur at extreme low mass ratios. No real time merger event has been documented, the only confirmed event, that of V1309 Sco, was only recognised in retrospect (Tylenda et al., 2011). Considerable advance has been made in defining likely instability parameters however finding systems that meet such criteria are rare (Wadhwa et al., 2024b). Although many hundreds of thousands of contact binary systems have been identified, only a small percentage are available for analysis without the need for large telescopes, sophisticated instruments and prohibitively long telescope times. Although methods have been developed to exclude likely stable systems based on survey photometry there is no definitive method to confirm potential orbital instability based purely on survey data. V717 And was selected from survey data because it demonstrated features of likely low mass ratio given the small amplitude and long total eclipse duration.
@@
Although the system is of extreme low mass ratio, orbital stability analysis suggests that the system is stable and not a potential merger candidate at this time. The presence of enhanced chromospheric activity although not a marker of orbital stability it can be a marker of enhanced magnetic activity. Enhanced magnetic activity may indicate significant magnetic stellar winds and potential enhanced angular momentum loss and period instability. The system, although not currently unstable, should at least undergo regular period monitoring.